OpenStudy (anonymous):

Solve the equation. 70 = –7(–2 – 2z)

OpenStudy (aroub):

70=14+14z 70-14=14z 56=14z 4=z

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

70=14+14z 70-14=14z 56=14z 56/14=z 4=z

OpenStudy (anonymous):

this would look more nice : \[70=14+14z\] \[70-14=14z\] \[56=14z\] \[\frac{56}{14}=z\] \[4=z \]